# Massless limit of massive self-interacting vector fields

### v2

@@ 1 Introduction
Pathology at Quantum level: Conventional approaches indicate that these theories do not have a smooth massless limit, with perturbative series becoming singular in mass, suggesting a violation of unitarity or failure of the regular perturbation theory Hell2022. This issue is not unique to vector fields; linearized massive gravity also suffers from a similar problem known as the van Dam-Veltman-Zakharov (vDVZ) discontinuity vanDam1970. This discontinuity arises because the longitudinal mode in massive gravity fails to decouple from the other degrees of freedom as the mass approaches zero. Both General Relativity and massive gravity are nonlinear theories that are typically solved using perturbation theory. However, there exists a scale known as the Vainshtein radius Vainshtein1972gravity, at which nonlinear terms become significant, causing perturbation theory to break down for longitudinal modes Vainshtein1972; Vainshtein1972gravity. This breakdown leads to a strong coupling regime in which the longitudinal mode decouples, restoring General Relativity’s predictions in the massless limit Vainshtein1972gravity. Recent work by Anamaria attempted to demonstrate a smooth massless limit for massive self-interacting theories using the Vainshtein mechanism Hell2022. However, our findings using the hyperbolic formulation suggest that the massless limit is not smooth at the classical level, in presence of quartic self coupling. We revisited the non-covariant formulation used by Anamaria and compared it with the hyperbolic formulation, which helps in determining the true behavior of the massless limit at the quantum level.
@@ 2 Hyperbolic Formulation of Nonlinear Proca Theory
In this section, we review the hyperbolic formulation for vector fields. We start with the Proca theory, which describes a massive vector field, and then proceed to self-interacting theories. We utilize the hyperbolic formulation to check the hyperbolicity of the equation of motion. The Lagrangian for the Proca theory is given by
${\cal{L}}_{PT}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{\mu^{2}}{2}X^{2}\ ,$ (1)

### v3

@@ 1 Introduction
Pathology at Quantum level: Conventional approaches indicate that these theories do not have a smooth massless limit, with perturbative series becoming singular in mass, suggesting a violation of unitarity or failure of the regular perturbation theory Hell2022. This issue is not unique to vector fields; linearized massive gravity also suffers from a similar problem known as the van Dam-Veltman-Zakharov (vDVZ) discontinuity vanDam1970. This discontinuity arises because the longitudinal mode in massive gravity fails to decouple from the other degrees of freedom as the mass approaches zero. Both General Relativity and massive gravity are nonlinear theories that are typically solved using perturbation theory. However, there exists a scale known as the Vainshtein radius Vainshtein1972gravity, at which nonlinear terms become significant, causing perturbation theory to break down for longitudinal modes Vainshtein1972; Vainshtein1972gravity. This breakdown leads to a strong coupling regime in which the longitudinal mode decouples, restoring General Relativity’s predictions in the massless limit Vainshtein1972gravity. Recent work attempted to demonstrate a smooth massless limit for massive self-interacting theories using the Vainshtein mechanism Hell2022. However, our findings using the hyperbolic formulation suggest that the massless limit is not smooth at the classical level, in presence of quartic self coupling. We revisited the non-covariant formulation and compared it with the hyperbolic formulation, which helps in determining the true behavior of the massless limit at the quantum level.
@@ 2 Hyperbolic Formulation of Nonlinear Proca Theory
The signature of the space-time metric is $(+,-,\dots,-)$.
In this section, we review the hyperbolic formulation for vector fields. We start with the Proca theory, which describes a massive vector field, and then proceed to self-interacting theories. We utilize the hyperbolic formulation to check the hyperbolicity of the equation of motion. The Lagrangian and the action for the Proca theory in a fixed space time with metric $g_{\mu\nu}$ is given by
${\cal{L}}_{PT}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{\mu^{2}}{2}X^{2}\ \quad\text{and,}\quad{\cal{S}}_{PT}=\int d^{4}x\sqrt{-g}\left(-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{\mu^{2}}{2}X^{2}\right)$ (1)